Factorise the following expressions.
q^2 –10q+21 ?????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

q^2–10q+21
Observed that, 21 = -7×-3 and -7+(-3) = -10
q^2–10q+21
= q^2–3q-7q+21
= q(q–3)–7(q–3)
= (q–7)(q–3)
This implies q^2–10q+21 = (q–7)(q–3)

Thanks